<html>
  <head>

    <meta http-equiv="content-type" content="text/html; charset=UTF-8">
  </head>
  <body>
    <p>Folks,</p>
    <p>Please find the notes <a moz-do-not-send="true"
href="https://halldweb.jlab.org/wiki-private/index.php/Hall_D_Staff_Meeting,_April_27,_2020">here</a>
      and below.</p>
    <p>  -- Mark</p>
    <p> </p>
    <div id="content" class="mw-body" role="main">
      <h1 id="firstHeading" class="firstHeading" lang="en"><span
          dir="auto">Hall D Staff Meeting, April 27, 2020</span></h1>
      <div id="bodyContent" class="mw-body-content">
        <div id="siteSub">From Gluex PRIVATE Wiki</div>
        <div id="mw-content-text" dir="ltr" class="mw-content-ltr"
          lang="en">
          <p>Present: Fernando Barbosa, Eugene Chudakov (chair), Mark
            Dalton, Alexandre Deur, Hovanes Egiyan, Sergey Furletov,
            Mark Ito, Igal Jaegle, Keigo Mizutani, Lubomir Pentchev,
            Patrizia Rossi, Elton Smith, Sascha Somov, Simon Taylor, Tim
            Whitlatch, Beni Zihlmann
          </p>
          <h2><span class="mw-headline" id="Announcements.2C_Eugene">Announcements,
              Eugene</span></h2>
          <ul>
            <li> Training will be offered for engineers</li>
            <li> Return to work being discussed, PPE requirements being
              discussed</li>
            <li> PAC has shifted to August 10-14. Proposals due June 22.</li>
            <li> New hires postponed. Alexander Austregesilo still with
              CMU.</li>
          </ul>
          <h2><span class="mw-headline" id="Announcements.2C_Patrizia">Announcements,
              Patrizia</span></h2>
          <ul>
            <li> EIC mini-review for CD1 will be held Sep.-Oct., full
              review in January</li>
          </ul>
          <h2><span class="mw-headline"
              id="Reports.2C_Other_Announcements.2C_Status_Updates">Reports,
              Other Announcements, Status Updates</span></h2>
          <ul>
            <li> Sergey: RAID system fixed</li>
            <li> Sascha: inspected crystals from SICCAS. They look good.
              Invoice can be paid and work on second batch started. Jane
              did good job with pre-selection in Shanghai</li>
            <li> Elton received 11 pages of comments on NIM paper. All
              text comments, save those that need discussion, dealt
              with. Turning to comments on figures.</li>
            <li> Beni reported that the monitoring launch on batch 4 of
              recent data complete. Batch 5 underway.</li>
            <li> Sascha: PrimEx processing almost ready. Change
              implemented to require both ADC and TDC hits for tagger
              counters.</li>
            <li> PAC proposals:
              <ul>
                <li> Alexandre: GDH proposal folks waiting on assessment
                  from reviewers. Discussion held on how to address
                  comments from meeting. Mark D. is streamlining the
                  proposal. Alexandre extending simulations to lower
                  energy range to overlap previous data. Discussing
                  impact with theorists.</li>
                <li> Elton: Rory is managing document for CPP jeopardy.
                  Writing assignments made. On NPP: message from Matt
                  Shepherd says we need proposal to collaboration by May
                  6. Group working to address comments from Barry
                  Ritchie on backgrounds. Will estimate relative level
                  of some backgrounds using PrimEx single π<sup>0</sup>
                  result.</li>
                <li> Eugene: Matt working on GlueX jeopardy.</li>
              </ul>
            </li>
            <li> Eugene: Justin asked about magnetic field accuracy and
              DIRC resolution. Eugene look at scheme to correct with
              multipoles tied to measurements in Hall
              <ul>
                <li> Simon: magnetic field had been excluded from DIRC
                  volume in MC. Fixed now. Still residual asymmetry
                  between charge species. Could be due to inaccuracies
                  in magnetic field map.</li>
                <li> Lubomir: including hit from GEM should improve
                  resolution</li>
              </ul>
            </li>
            <li> Eugene: constant term in calorimeter resolution quoted
              in NIM paper at 4%, rather high
              <ul>
                <li> Elton: for BCAL due to averaging over angles, for
                  FCAL: measurement over small energy range makes
                  separation of constant term from stochastic term
                  difficult</li>
              </ul>
            </li>
            <li> Sascha: SRC readiness review on May 7. Lubomir is PDL.
              Need to replace all active bases for CompCal.
              <ul>
                <li> Fernando discussed bases with Chris Cuevas. May
                  need new design to accommodate changes from Vladimir
                  Popov.</li>
              </ul>
            </li>
            <li> Tim on technical crew work:
              <ul>
                <li> Reviewing procedures, e.g., installing and removing
                  mirror boxes, tagger vacuum window.</li>
                <li> Work on DIRC electronics air cooling scheme.</li>
                <li> More work on planning for bringing up the detector</li>
                <li> Training</li>
              </ul>
            </li>
            <li> Fernando on electronics team activities:
              <ul>
                <li> Chris Stanislav working on boards for streaming
                  read-out of flash 125 boards</li>
                <li> Fernando in contact with CAEN on FCAL upgrade HV</li>
              </ul>
            </li>
            <li> Sergey: working on LDRD on FPGAs and event reduction
              for streaming read-out. Applications for EIC and Solid.
              Pre-proposal submitted. Full proposal due in May.
              <ul>
                <li> Beni asked about FPGA heat dissipation. FPGAs will
                  be in external crate connected by optical fiber to
                  data source. Cooling can be addressed with this
                  configuration.</li>
              </ul>
            </li>
            <li> Beni asked about schedule for return when/if Governor's
              stay-at-home order expires June 10. No clear answer yet.</li>
            <li> Sascha asked about summer students this year. No clear
              answer yet.</li>
          </ul>
        </div>
        <div class="printfooter"> Retrieved from "<a dir="ltr"
href="https://halldweb.jlab.org/wiki-private/index.php?title=Hall_D_Staff_Meeting,_April_27,_2020&oldid=10950">https://halldweb.jlab.org/wiki-private/index.php?title=Hall_D_Staff_Meeting,_April_27,_2020&oldid=10950</a>"
        </div>
      </div>
    </div>
    <div id="footer" role="contentinfo">
      <ul id="footer-info">
        <li id="footer-info-lastmod"> This page was last modified on 27
          April 2020, at 16:59.</li>
        <li id="footer-info-viewcount">This page has been accessed once.</li>
      </ul>
    </div>
  </body>
</html>